# Efficient and Accurate Spatial Mixing of Machine Learned Interatomic Potentials for Materials Science

### v1

@@ 4.6 Diffusion coefficients
To calculate the diffusion coefficients, individual defects were placed into periodic cells of size 16 $\times$ 16 $\times$ 16 unit cells ($\sim 8000$ atoms). Cells were heated to the target simulation temperature over 2 ps with a Langevin thermostat. For the Fe simulation, diffusion coefficients were measured at 800 K, 900 K, 1000 K and 1100 K. For W–He, diffusion coefficients were measured at 400 K, 600 K and 800 K. The reference simulations were then switched to NVE, whilst the ML/ML cells were switched to ’psuedo NVE’ with a weak Langevin thermostat (damping parameter - 2.0 ps) applied only to the cheap potential region. For the Fe simulation, each data point consisted of the average of five diffusion coefficients derived from the mean square displacements from non-overlapping time-lags of $5\times 1$ ns simulations. For the W–He simulations, each point was the average of twenty-five diffusion coefficients derived from the mean square displacements from non-overlapping time-lags from $25\times 60$ ps simulations. In all cases, the error was estimated from the standard deviation in the averaged diffusion coefficients. For the W–He ML/ML simulation, the seed atom was selected to be the single He present in the simulation, and regions were rebuilt around this seed atom every timestep. For the Fe ML/ML simulation, seed atoms were selected by finding the most highly coordinated atoms (on average) in the simulation every 100 timesteps. For this, measurements of atomic coordination $Z$ (with a cutoff of 2.0 Å) were taken every 10 timesteps, and these were averaged using fix ave/atom. New seed atoms were selected according to the criteria that $\langle Z\rangle>0.4$. Regions were rebuilt around these seed atoms every timestep. For the first 100 timesteps in the Fe ML/ML simulation, only the expensive potential was used. Both simulations used a core potential region of 6.0 Å around the defect atoms, as well as a blending region of 4.0 Å. For the Fe potential, a buffer width of 5.5 Å was used, whilst for the W–He potential, a width of 6.0 Å was used, matching the UF3 potential cutoff.
@@
All the data generated and analysed in the current study is available in the v0.1 version of the ML-MIX repository released on Zenodo [62].
@@
All underlying code for this study is available in the ML-MIX GitHub repository [57].
@@ Acknowledgements
We thank Christoph Ortner for helpful discussions. We also thank Lakshmi Shenoy and Matthew Nutter for providing the Fe and W-He ACE potential fitting scripts. FB is supported by a studentship funded by the UK Engineering and Physical Sciences Research Council supported Centre for Doctoral Training in Modelling of Heterogeneous Systems, Grant No. EP/S022848/1. JRK acknowledges funding from the NOMAD Centre of Excellence funded by the European Commission under grant agreement 951786. We acknowledge the University of Warwick Scientific Computing Research Technology Platform for assisting the research described within this study. Some of the calculations were performed using the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick. Sulis is funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium. TDS gratefully acknowledges support from ANR grants ANR-19-CE46-0006-1, ANR-23-CE46-0006-1, IDRIS allocation A0120913455 and an Emergence@INP grant from the CNRS.
@@
FB, TDS and JRK designed the research. FB developed the software, generated the data and performed the analyses. TDS and JRK provided supervision, guidance, and feedback throughout. FB wrote the paper, with input from TDS and JRK at all stages. All authors revised the paper and approved its final version.
Appendix A Expensive potential fitting parameters
Here we present the full set of potential fitting parameters for the Fe and W-He ACE potentials used in the study. The main fitting parameters are presented in table 6. Weights by config type for the Fe potential are shown in table 7. For the W–He ACE, we derive energy weights for configurations following the method suggested in [47] in which the energy weight on configuration $R$ is generated by
$w_{R{}}=\frac{1000}{\sqrt{N_{R}}}$ (6)
where $N_{R}$ is the number of atoms in configuration $R$. For each configuration, the force weights were set to 10.0 and the virial weights were set to 0.0. Virials were not used in the fitting in order to avoid any errors introduced from underconverged (insufficient $K$-point density) DFT, as suggested by the dataset author [63]. The exception to this were the single W atom configurations, which had an energy weight of 4000.0, a force weight of 100.0 and a virial weight of 200.0.
Table 6: Fe and W–He ACE potential parameters.
Fe ACE W–He ACE
Cutoff 5.5 5.0
Corr. order 3 3
Max. poly. deg. 20 20
Solver BLR BLR
Smoothness prior
strength
4 -
Table 7: Fe ACE configuration weighting.
Config type
Energy
weight
Force
weight
Virial
weight
default 30.0 1.0 1.0
slice_sample 80.0 0.1 1.0
prim_random 80.0 0.1 1.0
Appendix B Speedup and load balancing of mixed simulations
This section provides an investigation into the speedups for load balanced and non-load balanced ML-MIX simulations in silicon with a fixed expensive potential region size for (a) a fixed number of atoms and an increasing number of processors and (b) a fixed number of processors and an increasing number of atoms. It is shown that load balancing is most important when considering small domains ($<$30,000 atoms) and fewer MPI parallel processors ($<$15). It is shown that for small domains, the speedup variability between different numbers of processors is lessened by load balancing.
B.1 Method
MD simulations were conducted with Si domains at 0 K (frozen atoms). Two sets of simulations were run, set (a) where the number of atoms were fixed and the number of processors were increased between 1 and 48, and set (b) where the number of processors was fixed and the number of atoms were increased between $8\times 10^{3}$ and $10^{6}$. For each separate simulation domain, timing measurements were carried out with (i) only the 2_10 ACE Si cheap potential, (ii) only the 4_20 ACE Si expensive potential, (iii) load-balanced and (iv) non-load balanced ML/ML simulations containing both the expensive and cheap potentials. Within each simulation domain, the same number of timesteps was used for each timing measurement (i-iv), with the number selected such that the wall-time of the longest (the all-expensive) simulation was approximately 45 minutes. To build the expensive domain in each ML/ML simulation, single seed atoms were selected at the center of each cell, around which 6 Å core regions and 6 Å buffer regions were constructed. This meant that the number of expensive potential atoms was fixed across all ML/ML simulations, regardless of overall domain size.
For set (a), two atomistic domain sizes were chosen; a ‘small’ domain of 8000 atoms, and a ‘large’ domain of 262,144 atoms. For set (b), the number of processors was set at 27 for all domain sizes used. For each simulation domain investigated, the ‘upper-limit’ speedup $S_{\mathrm{UL}}$ was computed with
$S_{\mathrm{UL}}=\frac{NS_{\mathrm{C/E}}}{N_{\mathrm{E}}S_{\mathrm{C/E}}+N_{\mathrm{C}}},$ (7)
where $N$ is the total number of atoms in a simulation domain, $N_{\mathrm{E}}$ and $N_{\mathrm{C}}$ are the number of expensive and cheap atoms (including those in both buffer regions), and $S_{\mathrm{C/E}}$ is the measured speedup of the all-cheap simulation over the all-expensive simulation in the domain of interest.
Simulations were carried out on Dell PowerEdge C6420 compute nodes each with 2 x Intel Xeon Platinum 826 (Cascade Lake) 2.9 GHz 24-core processors; 48 cores per node; 192 GB DDR4-2933 RAM per node; 4 GB RAM per core.
B.2 Load balancing strategies
Load-balancing is performed via the fix balance command in LAMMPS. If the time keyword is specified LAMMPS attempts to dynamically decompose the overall domain into subdomains of equal computational cost. LAMMPS has two strategies to decompose domains - ‘brick’ and ‘tiled’. With the ‘brick’ strategy, each individual sub-domain is cuboidal, and is constrained to be joined at the corners. With the ‘tiled’ strategy, each domain is also cuboidal, but not necessarily constrained to connect at corners. For more details, please refer to the comm_style page of the LAMMPS documentation [64]. Within the fix balance command, two algorithms are available; the shift algorithm, compatible with both the ‘brick’ and ‘tiled’ decomposition strategies, and the rcb algorithm, compatible with just the ‘tiled’ decomposition strategy. For more detail on these algorithms, please refer to the fix balance page of the LAMMPS documentation [65]. Each load balancing strategy was investigated, and it was found for a simple domain with one central expensive region, the shift algorithm performed marginally better when domains were small and the rcb algorithm performed marginally better when domains were large. Throughout this paper, the shift algorithm is used with the brick domain decomposition strategy. We recommend users test each load-balancing strategy to determine which works best for their system.
B.3 Results and discussion
B.3.1 Fixed atoms, increasing processors
Figure 9: Speedup of ML/ML simulations over all-expensive reference simulations on an increasing number of processors for two domain sizes. In each case, the upper bound speedup (green line) is determined from equation 7. Left: Speedup on a small (8000 atom) domain. For this domain size, one can only attain near-upper bound speedups for small numbers of processors if a load-balancing strategy is used. The effect of load balancing is very pronounced, as without it all expensive atoms frequently end up on one sub-domain, creating simulation bottle-necks. Right: Speedup on a large (262,144 atom) domain. At this domain size, one attains a near upper-bound speedup for any number of MPI parallel domains, and load balancing has a relatively minor effect. All trend-lines are fitted with Gaussian process regression.
The results for simulation set (a) are displayed in Fig. 9. For the smaller (8000 atom) domain, it can be seen that with load balancing, a near-upper-bound speedup can be attained for a number of processors below $\sim$15. Above this, the load balanced simulation falls significantly below the upper-bound due to communication overheads taking up a significant fraction of the runtime. Without load balancing, the performance is on average significantly worse, but varies greatly between the number of processors chosen. Particularly poor performances are due to all of the expensive atoms landing on a single parallel domain. For the larger (262,144 atom) domain, it can be seen that near-upper bound speedup is approximately attained for any processor number, regardless of load-balancing. This arises due to every individual subdomain always being far larger than the expensive potential region.
B.3.2 Fixed processors, increasing atoms
Figure 10: The measured speedup of the mixed simulation over the all expensive simulation on 27 processors, for increasing numbers of atoms. The green line indicates the maximum possible speedup (with no overheads), computed from equation 7. It can be seen that for small numbers of atoms ($<$30,000) load balancing is important, and for large number of atoms ($\sim 10^{6}$) the speedup tends toward the upper bound. Trend lines have been fitted with Gaussian process regression.
Results for simulation set (b) in which the simulation domain size is increased at a fixed number of processors (27) is shown in Fig. 10. For smaller atomic domains ($<$30,000 atoms), it can be seen that load balancing is important, providing a considerable speedup. As the size of the domain gets larger, load-balancing becomes less important, and the overall speedup tends toward the overhead-free upper limit.
Appendix C Impact of constrained fitting
In this paper, the cheap potentials have been constrained to enforce elastic constant matching. In this section, we show that matching the elastic constants between our potentials leads to reduced errors in ML-MIX relaxed structures for less than 3% strain deviations from the bulk.
C.1 Methods
A periodic $10\times 10\times 10$ (8000 atom) bulk Si structure was constructed and sheared in the $xy$ to directions to strains between 0 and 7%. For each structure, a joint ML/ML relaxation was performed using ML-MIX. The central atom and the surrounding 6 Å were modeled expensively, with the remainder of the atoms were with the constrained or unconstrained Si 2_10 cheap potentials in two separate investigations. For the relaxation, a blending region of 6.0 Å was used, as well as a relatively loose tolerance of $\|\mathbf{F}\|_{2}<10^{-3}$ eV/Å. Smaller tolerances than this would frequently fail to converge. The strain error $Du$ was then found for each structure as
$Du=\|\nabla u-\nabla u_{\mathrm{ref}}\|_{2}$ (8)
where $u$ is the displacement field from the ML/ML relaxed structure and $u_{\mathrm{ref}}$ is the displacement field from the relaxed all-expensive reference.
C.2 Results and discussion
Figure 11: The strain error norm Du measured for sheared and relaxed bulk Si using an ML/ML simulation compared to the all-expensive simulation for different applied shear strain. The ML/ML simulation contains a central 6 Å sphere of the expensive potential surrounded by the cheap potential. $Du$ is far greater for small strains when using the unconstrained cheap potential, as the mismatched elastic constants to the expensive lead to Eshelby-inclusion type strain fields.
Fig. 11 shows the strain error in the ML/ML relaxed structures. The spatial distribution of the error in the strain fields matches that expected from an Eshelby inclusion, and it can be seen that for very small strains $\epsilon_{xy}<0.005$, the constrained potential exhibits very little error in the strain whilst the constrained potential error increases immediately. The constrained potential continues to have lower strain error up until $\epsilon_{xy}=0.03$.
Appendix D Energy conservation
A fundamental limitation of force mixing is a lack of energy conservation. Whilst this has been noted in previously [35], the features in a force-mixed simulation that may worsen or improve energy conservation are unclear. Here we present a brief investigation where we test the impact on energy conservation of (i) having a larger interface region between potentials, (ii) having a more flexible cheap potential and (iii) introducing a larger blending region between potentials. Whilst findings presented here are derived from ML/ML simulations using ML-MIX, the findings are applicable to force-mixing simulations more generally, including QM/MM simulations.
D.1 Methods
A periodic cube of Si was constructed which matches that described in section 2.3. The spherical central region of the domain (around the fixed stretched bond) was simulated expensively, with the remainder of the domain modeled cheaply. The expensive potential was the same Si ACE used in section 2.3. Three different cheap ACE potentials were used, fit according to the constrained fitting procedure outlined in section 4.1. The key parameters of these potentials are displayed in table 8. To investigate energy drift in the simulation set-up, a 50 ps, 300 K ML-MIX LAMMPS MD simulation was performed with a timestep of 1 fs. As the total energy of a force-mixed simulation is not defined, the energy was measured through evaluating snapshots output every ps with just the expensive potential.
Table 8: Si cheap potential parameters.
2_10 2_15 3_15
Cutoff 6.0 6.0 6.0
Corr. order 2 2 3
Max. poly. deg. 10 15 15
Two separate investigations were conducted for each cheap potential. (a) For a set-up with no blending region (abrupt force-mixing) the size of the spherical expensive potential region was varied between 4 and 10 Å, and (b) for a fixed 10 Å core + blending region, the amount of blending region is varied between 4 and 6 Å.
D.2 Results and discussion
D.2.1 Size of potential-potential interface
Figure 12: Energy drift over time for an abruptly mixed ML/ML simulation of bulk Si, using the 2_10 cheap potential. The size of the expensive region is varied between 4 and 10 Å. Left - the raw energy change for each simulation, showing that a larger expensive potential region causes a greater rate of energy increase. Right - energy change normalised by spherical interfacial contact area between potentials, showing that for a given two potentials in abrupt contact the rate of energy change per unit of contact area (energy flux) is constant.
The energy over time for the abruptly force-mixed ML/ML simulations are shown on the left of Fig. 12. Two facts are discernible from these plots: the energy drifts are approximately linear, and larger expensive potential regions lead to greater energy gain. The right hand side of Fig. 12 shows the energy gains normalised by the (approximately spherical) surface interface area between potentials. Under this normalisation, the curves collapse onto a single straight line, giving us the result that for two given potentials and a given blending regime the rate of energy increase per unit surface contact area is approximately constant. The gradient of this line can be thought of as the energy flux from the force mixing contact surface into the material caused by the local violation of Newton’s 3rd law.
D.2.2 Flexibility of cheap potential
Figure 13: Left - energy flux for abruptly mixed ML/ML simulations with different cheap potentials of increasing complexity. More complex cheap potentials lead to orders of magnitude lower energy drift. Right - energy flux for ML/ML simulations with the 2_10 potential with different blending widths. The blue bar corresponds to abrupt mixing, whilst the other bars correspond to increasing amounts of blending. Increasing the blending width decreases the energy flux from the potential contact surface.
Energy fluxes from ML/ML simulations of three different cheap potentials are shown on the left-hand side of Fig. 13. Note that as the cheap potential complexity increases, the energy flux from the internal potential/potential contact surface decreases by orders of magnitude. The rationale for this is straightforward; more complex cheap potentials with greater numbers of basis functions have the flexibility to better approximate the local potential energy surface of the expensive potential, causing the violations of Newton’s 3rd law to be lesser in magnitude.
D.2.3 Width of blending region
The right-hand side of Fig 13 shows how adding a blending region in the ML-MIX simulation can decrease the energy flux. The blue bar corresponds to abrupt mixing, and it can be seen that as core region is swapped out for blending region the amount of energy drift in the simulation decreases.

### v2

@@ 4.8 Diffusion coefficients
To calculate the diffusion coefficients, individual defects were placed into periodic cells of size 16 $\times$ 16 $\times$ 16 unit cells ($\sim 8000$ atoms). Cells were heated to the target simulation temperature over 2 ps with a Langevin thermostat. For the Fe simulation, diffusion coefficients were measured at 800 K, 900 K, 1000 K and 1100 K. For W–He, diffusion coefficients were measured at 400 K, 600 K and 800 K. The reference simulations were then switched to NVE, whilst the ML/ML simulations had a weak Langevin thermostat (damping parameter - 2.0 ps) applied only to the cheap potential region. For the Fe simulation, each data point consisted of the average of five diffusion coefficients derived from the mean square displacements from non-overlapping time-lags of $5\times 1$ ns simulations. For the W–He simulations, each point was the average of twenty-five diffusion coefficients derived from the mean square displacements from non-overlapping time-lags from $25\times 60$ ps simulations. In all cases, the error was estimated from the standard error in the averaged diffusion coefficients. For the W–He ML/ML simulation, the seed atom was selected to be the single He present in the simulation, and regions were rebuilt around this seed atom every timestep. For the Fe ML/ML simulation, seed atoms were selected by finding the most highly coordinated atoms (on average) in the simulation every 100 timesteps. For this, measurements of atomic coordination $Z$ (with a cutoff of 2.0 Å) were taken every 10 timesteps, and these were averaged using fix ave/atom. New seed atoms were selected according to the criteria that $\langle Z\rangle>0.4$. Regions were rebuilt around these seed atoms every timestep. For the first 100 timesteps in the Fe ML/ML simulation, only the expensive potential was used. Both simulations used a core potential region of 6.0 Å around the defect atoms, as well as a blending region of 4.0 Å. For the Fe potential, a buffer width of 5.5 Å was used, whilst for the W–He potential, a width of 6.0 Å was used, matching the UF3 potential cutoff.
4.9 W dislocation glide
Simulating the thermally activated glide of W screw dislocations is known to require $\approx 10^{5}$ atoms due to finite-size effects [cereceda_2012, allera_2024, Bacon_2009]. These are (i) in $x$, the cell length must be large enough to prevent self-interaction and to ensure that the work done can dissipate without substantial heating, (ii) in $y$, the surface must be far enough away to prevent attraction and dislocation annihilation, and (iii) in $z$, to ensure that the dislocation moves by the correct mechanism of the nucleation and propagation of kink pairs. In this work, the dimensions of the simulation cell in the glide direction (along $x=[10\overline{1}]$), glide plane normal (along $y=[\overline{1}2\overline{1}]$), and dislocation line (along $z=[111]$) were chosen to be $223~\mathrm{\AA }\times 221~\mathrm{\AA }\times 68~\mathrm{\AA }\ (25~\textit{b}$), respectively, for a total of 195,000 atoms. At each temperature simulated, the thermally expanded lattice parameter was measured and the cell was rescaled to match. In each simulation, the dislocation was placed at the centre of the simulation cell, where the atomic displacement field was generated using the matscipy dislocation module [matscipy]. Periodicity was imposed in $x$ and $z$, with free surfaces in $y$. Shear stresses $\sigma_{\mathrm{yz}}$ were applied to these free surfaces, simulating {112} loading. During glide, the dislocation moved on alternating {110} planes.
Using a 2 fs timestep, the cell was equilibrated to the chosen temperature for 50 ps with a Langevin thermostat. Following this, a 250 ps production run was conducted, from which dislocation positions were extracted (using OVITO dislocation analysis [Stukowski_2010]) and an average dislocation velocity was calculated. In the all-expensive simulations, an NVE ensemble was used, whilst in the mixed simulations, a weak Nosé-Hoover thermostat was applied to the whole system (damping time 10 ps) in order to counteract the gradual heating that occurred due to energy flux from the ML/ML boundary. A short initial test was conducted to ensure that this thermostat did not impact dynamics.
The dislocation core was identified as the expensive region using Common Neighbor Analysis (CNA) in LAMMPS. The radius of the core, blending and buffer regions were 5.0, 4.0 and 5.0 $\mathrm{\AA }$, respectively. The expensive region corresponded to $\approx 1\%$ of the total atoms in the simulation cell.
4.10 He implantation into W
Table 6: The cell size and number of atoms used for He implantation simulations at energies up to 100 eV.
He Energy (eV) Cell Size (Å) Number of Atoms
5-30 $\times39.8$$\times$46.1 5070
35-40 $\times46.1$$\times$62.0 9000
50-60 $\times46.1$$\times$93.8 13500
70 $\times46.1$$\times$125.6 18000
80 $\times46.1$$\times$157.4 22500
90 $\times46.1$$\times$189.2 27000
100 $\times46.1$$\times$221.0 31500
[…]
To generate the initial configurations for each implantation, a He atom was placed in a uniformly sampled random position 5.0 Å above the W surface and given a velocity in z-z corresponding to the investigated kinetic energy (between 5 and 100 eV). An initial simulation timestep was selected corresponding to an initial He per-timestep displacement of 0.02 Å. The initial surface impact simulation was then run for 750 timesteps, which would correspond to a total straight-line He distance of 15.0 Å. At the end of this, the He atom position was measured. If it was found to be more than 2.0 Å above the surface, and with a positive $z$ velocity, the simulation was stopped with the result being that the atom reflected. If not, the timestep was updated to either 0.02 Å divided by average He velocity or 0.1 fs, whichever was smaller, and a further 0.05 ps was run. This was repeated until either the He escaped the surface or a total time of 1 ps passed. For all energies, 1 ps was enough time for the He atom to thermalise.
All expensive simulations were carried out with He energies up to 40 eV. For each energy, 4 sets of 25 implantation attempts were carried out, with each set of 25 sharing the same thermalized slab structure.
For the mixed simulations, the expensive region was set to track the He atom, with a core radius of 6.0 Å a blending radius of 4.0 Å and a buffer radius of 8.0 Å. The thermalization was entirely conducted using the cheap potential. For all energies, 5 sets of 100 repeats were used.
The potentials were evaluated in LAMMPS on 40 GB A100 GPUs using the Kokkos [kokkos1, kokkos2] accelerated variants of the symmetrix package for MACE [Kovacs_2025_maceoff] and ML_SNAP package for SNAP [Thompson_2015]. In order to conduct mixed simulations where both the expensive and cheap potentials were evaluated on the same GPU, a Kokkos version of ML-MIX was developed. This can be found in v0.3 of ML-MIX onwards, and allows for force-mixing on GPU of many kokkos-compatible LAMMPS pair_styles.
@@ 4.10 He implantation into W
Data generated and analysed in the current study is available in the v0.3 version of the ML-MIX repository released on Zenodo [ml-mix-zenodo].
@@
The ML-MIX package along with scripts necessary to reproduce the results presented here are available in the ML-MIX GitHub repository [ML-MIX].
@@ Acknowledgements
We thank Christoph Ortner and Albert P. Bartók for helpful discussions. We also thank Lakshmi Shenoy for the Fe fitting script. FB and MN are supported by studentships funded by the UK Engineering and Physical Sciences Research Council supported Centre for Doctoral Training in Modelling of Heterogeneous Systems, Grant No. EP/S022848/1. MN is also supported by AWE Nuclear Security Techonologies. We acknowledge the University of Warwick Scientific Computing Research Technology Platform for assisting the research described within this study. Some of the calculations were performed using the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick. Sulis is funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium. We are grateful for computational support from the UK national high performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and funded by EPSRC Grant No. EP/X035891/1. TDS gratefully acknowledges support from ANR grants ANR-19-CE46-0006-1, ANR-23-CE46-0006-1, IDRIS allocation A0120913455 and an Emergence@INP grant from the CNRS.
@@
FB, MN, TDS and JRK designed the research. FB developed the ML-MIX software, generated the data and performed the analyses for all examples except the W screw dislocation glide, which was conducted by MN. TDS and JRK provided supervision, guidance, and feedback throughout. FB wrote the paper, with input from TDS and JRK at all stages, as well as MN at the final stage. A draft of the W screw dislocation sections was provided by MN. All authors revised the paper and approved the final version.